# Lorentzian path-integral of Robin Universe

### v1

@@ Abstract
The gravitational path-integral of Gauss-Bonnet gravity is investigated and the transition from one spacelike boundary configuration to another is analyzed. Of particular interest is the case of Neumann and Robin boundary conditions which is known to lead to stable Universe in Einstein-Hilbert gravity in four spacetime dimensions. After setting up the variational problem and computing the necessary boundary terms, the transition amplitude can be exactly computed in the mini-superspace approximation. The $\hbar\to 0$ limit brings out the dominant pieces in the path-integral which is traced to an initial configuration corresponding to Hartle-Hawking no-boundary Universe. A deeper study involving Picard-Lefschetz methods not only allow us to find the integration contour along which the path-integral becomes convergent but also aids in understanding the crossover from Euclidean to Lorentzian signature. Saddle analysis further highlights the boundary configurations giving dominant contribution to the path-integral which is seen to be those corresponding to Hartle-Hawking no-boundary proposal and agrees with the exact computation. To ensure completeness, a comparison with the results from Wheeler-DeWitt equation is done.
@@ I Introduction
In quantum field theory, path integral is a useful tool for computing the transition amplitude from one initial configuration to a final one. It allows us to obtain information about physical systems, most importantly scattering cross sections/S-matrix. In flat spacetime quantum field theory, the path integral computation involves the integration over the fields defined on the background of fixed flat spacetime, obeying appropriate boundary conditions.
However, when background spacetime is curved and/or is dynamically evolving as in the case of an expanding universe, defining a path-integral over fields is technically more involved. One has to carefully tackle issues like renormalizability, gauge-fixing, regularization, measure, boundary conditions and integration contour. Dealing with the last one comes with additional complications, which we aim to discuss here.
The gravitational path-integral (ignoring matter sector) with metric as the only degree of freedom can be schematically written as
@@
where $g_{\mu\nu}$ is the metric, $S[g_{\mu\nu}]$ is the corresponding action, $R$ is the Ricci scalar, $G_{N}$ is the Newton’s gravitational constant, $\Lambda$ is the cosmological constant , ()($\cdots$)( ⋯ ) are the higher derivative curvature terms. The path integral is defined on manifold $\mathcal{M}$ with boundary $\partial\mathcal{M}$. $\rm Bd_{i}$ and $\rm Bd_{f}$ refer to field configurations at the initial and final boundaries, respectively.
@@
It has two unknown time-dependent functions: lapse $N_{p}(t_{p})$ and scale-factor $a(t_{p})$, $k=(0,\pm 1)$ is the curvature, and ${\rm d}\Omega_{D-2}$ is the metric for the unit sphere in $D-2$ spatial dimensions. In this reduced space, although gravitational waves don’t exist, however diffeomorphism invariance still shows up in the time co-ordinate $t_{p}$ and the dynamical scale-factor $a(t_{p})$. This reduced set-up is enough to explore the effects of boundary conditions and integration contours in the gravitational path-integral.
The Feynman path integral in this reduced gravitational phase space is
@@
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}N_{p}{\cal D}\pi{\cal D}a(t_{p}){\cal D}p{\cal D}{\cal C}{\cal D}\bar{P}\exp\biggl{[}\frac{i}{\hbar}\int_{0}^{1}{\rm d}t_{p}\left(N^{\prime}_{p}\pi+a^{\prime}p+{\cal C}^{\prime}\bar{P}-N_{p}H\right)\biggr{]}\,,$ (4)
@@
where along with the integration over the scale-factor $a(t_{p})$, lapse $N_{p}(t_{p})$ and the Fermionic ghost ${\cal C}$, we also have an integration over their respective conjugate momenta given by $p$, $\pi$, and $\bar{P}$. It should be specified that the $({}^{\prime})$ here denotes derivative with respect to $t_{p}$, and without any loss of generality the time co-ordinate is chosen to range from $0\leq t_{p}\leq 1$. ${\rm Bd}_{i}$ and ${\rm Bd}_{f}$ refers to the field configuration at initial ($t_{p}=0$) and final ($t_{p}=1$) boundaries respectively. The Hamiltonian constraint $H$ consists of two parts: gravitational part and ghost
$H=H_{\rm grav}[a,p]+H_{\rm gh}[N,\pi,{\cal C},\bar{P}]\,,$ (5)
@@
where $H_{\rm grav}$ is the Hamiltonian for the gravitational action and the Batalin-Fradkin-Vilkovisky (BFV) Batalin:1977pb ghost Hamiltonian is denoted by $H_{\rm gh}$ 111 The BFV ghost generalizes the standard Fadeev-Popov ghost based on BRST symmetry. Usual gauge theories constraint algebra forms a Lie algebra. However, in the case of gravitational theories respecting diffeomorphism invariance the constraint algebra doesn’t close, thereby requiring BFV quantization. In the reduced phase-space of mini-superspace approximation although there is only one constraint (Hamiltonian $H$) thereby leading to trivial closure of algebra, but still, BFV quantization is preferable.. Even though the phase space is quite reduced in the mini-superspace approximation, there is still time reparametrization invariance which need to be suitably gauge-fixed. A simple gauge choice is the ADM gauge also known as proper-time gauge $N^{\prime}_{p}=0$. More details on BFV quantization process and BFV ghost can be found in the seminal works Teitelboim:1981ua ; Teitelboim:1983fk ; Halliwell:1988wc .
@@
A large part of the above gravitational path-integral mentioned in eq. (4) can be performed exactly. This includes path-integral over the ghost-sector: ${\cal C}$ and $\bar{P}$; the conjugate momenta: $\pi$ and $p$. The path-integral left behind after this is
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{0^{+}}^{\infty}{\rm d}N_{p}\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}a(t_{p})\,\,e^{iS[a,N_{p}]/\hbar}\,.$ (6)
@@
This path-integral is easy to interpret: $\int{\cal D}a(t_{p})\,\,e^{iS[a,N_{p}]/\hbar}$ gives the amplitude of transition for the Universe to evolve from one boundary configuration to another in the proper time $N_{p}$, while for the lapse integration one needs to consider paths of every proper duration $0<N_{p}<\infty$. This choice leads to causal evolution from one boundary configuration ${\rm Bd}_{i}$ to another ${\rm Bd}_{f}$ as discussed in Teitelboim:1983fh , where $a_{0}<a_{1}$ will imply expanding Universe while $a_{0}>a_{1}$ will imply contracting Universe.
@@
The gravitational action in eq. (7) falls in the class of lovelock gravity theories Lovelock:1971yv ; Lovelock:1972vz ; Lanczos:1938sf , and is a special type of higher-derivative gravity consisting of only two-time derivatives of metric-field. This has the privilege of keeping the dynamical evolution equation of the metric field second order in time. It should be pointed out that Gauss-Bonnet gravitational theories also arises in the low-energy effective action of the heterotic string theory with $\alpha>0$ Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx . Interestingly, this coupling also has received observational constraints Chakravarti:2022zeq coming from the study of the gravitational wave (GW) data of the event GW150914, which also offered the first observational confirmation of the area theorem Isi:2020tac .
@@
In this short review article, we will focus on the path integral given in the eq. (6) for the gravitational action specified by eq. (7), and discuss the effects of the Neumann/Robin boundary conditions (NBC/RBC) on the transition amplitude (see York:1986lje ; Brown:1992bq ; Krishnan:2016mcj ; Witten:2018lgb ; Krishnan:2017bte for the role played by boundary terms). Imposing RBC at the initial boundary is like specifying a combination of scale factor and its conjugate momenta to be fixed. A particular class of the same is “linear” combination of scale factor and its conjugate momenta, characterized by a single parameter. These RBCs interpolate between Dirichlet BC and Neumann BC. Past studies involving specifying DBCs at the initial boundary showed unsuppressed growth of fluctuations around the saddles leading to instability in the no-boundary proposal of the Universe Feldbrugge:2017kzv ; DiTucci:2018fdg ; Lehners:2018eeo ; Feldbrugge:2017fcc ; Feldbrugge:2017mbc ; Matsui:2024bfn (see Lehners:2023yrj for review on the no-boundary proposal). These studies motivated us to investigate whether NBC or RBC can lead to stable behavior of perturbations, also at the same time highlighting that Dirichlet BC is perhaps not the right boundary condition for gravity Krishnan:2016mcj ; Krishnan:2017bte .
@@
Explorations with the Neumann boundary condition (NBC) have shown that gravitational fluctuations are well-behaved in the no-boundary proposal of the Universe DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; Narain:2022msz . In the case of mini-superspace approximation with metric having ×$\mathbb{R}\times\mathbb{S}^{3}$ it is possible to do the computation of the path-integral exactly Narain:2022msz . These exact computation showed some interesting features where it was seen that our Universe undergoes signature changes from Euclidean to Lorentzian Narain:2022msz . Studies involving RBCs with Einstein-Hilbert gravity were done in DiTucci:2019dji ; DiTucci:2019bui in the WKB approximation using Picard-Lefschetz methods. In all these works it was noticed that perturbations were well-behaved. Exact computations imposing RBCs at the initial boundary have been done in Ailiga:2023wzl . Effects of Gauss-Bonnet addition has been throughly investigated in Narain:2021bff ; Narain:2022msz ; Ailiga:2023wzl .
@@
This review also focuses on the impact of having Gauss-Bonnet addition in the gravitational action and the role it plays at the boundaries. This is crucial as the path-integral is sensitive to the boundary conditions. Retaining the consistency of the variational problem, the boundary condition is appropriately chosen. This still doesn’t guarantee that the solutions respecting these will be a stable configuration. But it is crucial to work those boundary conditions which lead to stable perturbations. This eventually means that in path-integral we only sum over stable geometries.
@@
Such path-integrals need to be analyzed appropriately in a systematic manner in the arena of complex analysis. Recent work Kontsevich:2021dmb on the ‘allowability’ criterion proposes a simple method to determine the allowable complex metric on which quantum field theories can be consistently defined. It is however yet to be understood and established whether this proposal is necessary or sufficient (see also recent works Witten:2021nzp and Lehners:2021mah ). Our approach avoids the standard process of Wick-rotation and aims to directly tackle the Lorentzian gravitational path-integral using the Picard-Lefschetz (PL) method.
@@
Picard-Lefschetz (PL) methodology generalizes the notion of standard Wick-rotation by suitably changing it to handle highly oscillatory integrals in a reliable manner. It provides a framework that considers all the saddles of the path-integral and systematically picks only the ones which eventually lead to convergent path-integral. The original integration contour gets appropriately deformed into the integrals over the steepest descent contours passing through such saddles. Such contours termed Lefschetz thimbles together form the generalized Wick-rotated contour. In the Lorentzian quantum cosmology studies, these methods were utilized to understand the nature of the path integral Feldbrugge:2017kzv ; Feldbrugge:2017fcc ; Feldbrugge:2017mbc ; Vilenkin:2018dch ; Vilenkin:2018oja ; Rajeev:2021xit and the effects of the boundary conditions DiTucci:2019dji ; DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; Narain:2022msz ; Ailiga:2023wzl 222 Earlier studies involving complex analysis methods to understand Euclidean gravitational path integral was done in Hawking:1981gb ; Hartle:1983ai . Boundary effects studies in the Euclidean quantum cosmology was done in tunnelling proposal made in Vilenkin:1982de ; Vilenkin:1983xq ; Vilenkin:1984wp and no-boundary proposal made in Hawking:1981gb ; Hartle:1983ai ; Hawking:1983hj . Later works investigating the contour choice in the path-integral involving the usage of complex analysis methods were also done in Halliwell:1988ik ; Halliwell:1989dy ; Halliwell:1990qr . .
@@
The plan of this article is as follows: In section I, we present the introduction and discuss the motivation for studying the gravitational path integral. Section II discusses about the mini-superspace approximation and outlines the surface terms required for the consistent variational problem. In section III, we compute the exact transition amplitude for Einstein-Gauss-Bonnet gravity with Neumann/Robin boundary conditions and analyze the semiclassical behaviour in the section IV. In section IV.3 we discusses the significance of the Robin boundary condition. A brief overview of the Picard-Lefschetz technique is provided in the section V, followed by its application in the section VI to compute the saddle point results and verify their consistency with the results obtained from the exact analysis. Section VII outlines the connection between the path integral and canonical approaches. Finally, the article concludes with a summary and outlook in Section VIII.

### v2

# Lorentzian Robin Universe of Gauss-Bonnet Gravity

@@ Abstract
The gravitational path-integral of Gauss-Bonnet gravity is investigated and the transition from one spacelike boundary configuration to another is analyzed. Of particular interest is the case of Neumann and Robin boundary conditions which is known to lead to a stable Universe in Einstein-Hilbert gravity in four spacetime dimensions. After setting up the variational problem and computing the necessary boundary terms, the transition amplitude is computed exactly in the mini-superspace approximation. The $\hbar\to 0$ limit brings out the dominant pieces in the path-integral which is traced to an initial configuration corresponding to Hartle-Hawking no-boundary Universe. A deeper study involving Picard-Lefschetz methods not only allow us to find the integration contour along which the path-integral becomes convergent but also aids in understanding the crossover from Euclidean to Lorentzian signature. Saddle analysis further highlights the boundary configurations giving dominant contribution to the path-integral which is seen to be those corresponding to Hartle-Hawking no-boundary proposal and agrees with the exact computation. To ensure completeness, a comparison with the results from Wheeler-DeWitt equation is done.
KEYWORDS: Lorentzian Path-Integral, Boundary Conditions, Gauss-Bonnet Gravity, Picard-Lefschetz Methods, No-Boundary Universe, Wheeler-DeWitt Equation.
@@ I Introduction
In quantum field theory, path integral is a useful tool for computing the transition amplitude from one initial configuration to a final one. It allows us to obtain physical information about the system, most importantly scattering cross sections/S-matrix. In flat/curved spacetime quantum field theory, the path integral computation involves the integration over the fields defined on the background of fixed flat/curved spacetime, obeying appropriate boundary conditions.
However, when gravity is present and/or background spacetime is dynamically evolving, as in the case of an expanding universe, defining a path-integral over fields is technically more involved. One has to carefully tackle issues like renormalizability, gauge-fixing, regularization, measure, boundary conditions and integration contour. Dealing with the last one comes with additional complications, which we aim to discuss here.
The pure gravitational path-integral (ignoring matter sector) with metric as the only degree of freedom can be schematically written as
@@
where $g_{\mu\nu}$ is the metric, $S[g_{\mu\nu}]$ is the corresponding gravity action, $R$ is the Ricci scalar, $G_{N}$ is Newton’s gravitational constant, $\Lambda$ is the cosmological constant, ()($\cdots$)( ⋯ ) includes the higher derivative curvature terms. The path integral is defined on manifold $\mathcal{M}$ with boundary $\partial\mathcal{M}$. $\rm Bd_{i}$ and $\rm Bd_{f}$ refer to field configurations at the initial and final boundaries, respectively.
@@
It has two unknown time-dependent functions: lapse $N_{p}(t_{p})$ and scale-factor $a(t_{p})$, $k=(0,\pm 1)$ is the scalar curvature, and ${\rm d}\Omega_{D-2}$ is the metric for the unit sphere in $D-2$ spatial dimensions. In this reduced space, although gravitational waves don’t exist, a reduced version of diffeomorphism invariance still shows in the time co-ordinate $t_{p}$ and the dynamical scale-factor $a(t_{p})$. This reduced setup is enough to explore the effects of boundary conditions and integration contour in the gravitational path-integral.
The path integral in this reduced gravitational phase space is
$G[{\rm Bd}_{f},{\rm Bd}_{i}]$ $=\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}N_{p}{\cal D}\pi{\cal D}a(t_{p}){\cal D}p{\cal D}c{\cal D}\bar{c}{\cal D}\rho{\cal D}\bar{\rho}$ (4)
@@
×$\quad\quad\times\exp\biggl{[}\frac{i}{\hbar}\int_{0}^{1}{\rm d}t_{p}\left(N^{\prime}_{p}\pi+a^{\prime}p+c^{\prime}\rho+\bar{c}^{\prime}\bar{\rho}-N_{p}H\right)\biggr{]}\,,$
@@
where along with the integration over the scale-factor $a(t_{p})$, lapse $N_{p}(t_{p})$ the Fermionic ghost $c$ and the anti-ghost $\bar{c}$, we also have an integration over their respective conjugate momenta given by $p$, $\pi$, $\rho$, and $\bar{\rho}$ It should be specified that the $({}^{\prime})$ here denotes derivative with respect to $t_{p}$, and without any loss of generality the time co-ordinate is chosen to range from $0\leq t_{p}\leq 1$. ${\rm Bd}_{i}$ and ${\rm Bd}_{f}$ refers to the field configuration at initial ($t_{p}=0$) and final ($t_{p}=1$) boundaries respectively. The Hamiltonian constraint $H$ consists of two parts: gravitational part and ghost
$H=H_{\rm grav}[a,p]+H_{\rm gh}[N,\pi,c,\bar{c},\rho,\bar{\rho}]\,,$ (5)
@@
where $H_{\rm grav}$ is the Hamiltonian for the gravitational action and the Batalin-Fradkin-Vilkovisky (BFV) Batalin:1977pb ghost Hamiltonian is denoted by $H_{\rm gh}$. BFV ghost generalizes the standard Fadeev-Popov ghost based on BRST symmetry. Unlike the gauge theories, in gravitational theories, the constraint algebra doesn’t close, thereby requiring BFV quantization. In minisuperspace, where we have only one constraint (Hamiltonian Constraint), the algebra trivially closes, but still, BFV quantization is preferable. Even though the phase space is quite reduced in the mini-superspace approximation, there is still time reparametrization invariance, which needs to be suitably gauge-fixed. A simple gauge choice is the ADM gauge also known as proper-time gauge $N^{\prime}_{p}=0$. More details on the BFV quantization process and BFV ghost can be found in the seminal works Teitelboim:1981ua ; Teitelboim:1983fk ; Halliwell:1988wc .
@@
A large part of the above gravitational path-integral mentioned in eq. (4) can be performed exactly. This includes path-integral over the ghost-sector: $c,\bar{c}$ and $\rho,\bar{\rho}$; the conjugate momenta: $\pi$ and $p$. The path integral over the ghosts and the corresponding momenta leads to a lapse-independent constant term. Ignoring this constant term, the path-integral left behind is
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{\bf C}{\rm d}N_{p}\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}a(t_{p})\,\,e^{iS[a,N_{p}]/\hbar}\,.$ (6)
@@
This path-integral can be interpreted as follows: $\int{\cal D}a(t_{p})\,\,e^{iS[a,N_{p}]/\hbar}$ gives the transition amplitude for the Universe to evolve from one boundary configuration to another in the proper time $N_{p}$, where ${\bf C}$ refers to the contour of Lapse integration. Depending on the choice of integration range ${\bf C}$ takes, eq. (6) leads to different objects. The $0<N_{p}<\infty$ choice leads to causal evolution from one boundary configuration ${\rm Bd}_{i}$ to another ${\rm Bd}_{f}$ as discussed in Teitelboim:1983fh , where $a_{0}<a_{1}$ will imply expanding Universe while $a_{0}>a_{1}$ will imply contracting Universe. Whereas the other choice, $-\infty<N_{p}<\infty$, takes the gauge invariance into account. For $-\infty<N_{p}<0$, $a_{0}<a_{1}$ will imply an expanding Universe while $a_{0}>a_{1}$ will imply contracting Universe where the subscript is with respect to $t_{p}$ coordinate. The $0<N_{p}<\infty$ leads to the (causal) propagator, while $-\infty<N_{p}<\infty$ leads to the wavefunction of the universe, which is a solution to the Wheeler-DeWitt equation. In this discussion, we focus on the latter case, with the lapse integration range $-\infty<N_{p}<\infty$. This choice allows us to account for all the dominant saddles in the path integral.
@@
The gravitational action in eq. (7) falls in the class of lovelock gravity theories Lovelock:1971yv ; Lovelock:1972vz ; Lanczos:1938sf , and is a special type of higher-derivative gravity consisting of only two-time derivatives of metric-field. This has the privilege of keeping the dynamical evolution equation of the metric field second order in time. It should be pointed out that Gauss-Bonnet gravitational theories also arises in the low-energy effective action of the heterotic string theory with $\alpha>0$ Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx . Interestingly, this coupling also has received observational constraints Chakravarti:2022zeq coming from the study of the gravitational wave (GW) data of the event GW150914, which also offered the first observational confirmation of the area theorem Isi:2020tac .
@@
In this short review article, we will focus on the path integral given in the eq. (6) for the gravitational action specified by eq. (7), and discuss the effects of the Neumann/Robin boundary conditions (NBC/RBC) on the transition amplitude (see York:1986lje ; Brown:1992bq ; Krishnan:2016mcj ; Witten:2018lgb ; Krishnan:2017bte for the role played by boundary terms). Imposing RBC at the initial boundary is like specifying a combination of scale factor and its conjugate momenta to be fixed. A particular class of the same is “linear” combination of scale factor and its conjugate momenta, characterized by a single parameter. These RBCs interpolate between Dirichlet BC and Neumann BC. Past studies involving specifying DBCs at the initial boundary showed unsuppressed growth of fluctuations around the saddles leading to instability in the no-boundary proposal of the Universe Feldbrugge:2017kzv ; DiTucci:2018fdg ; Lehners:2018eeo ; Feldbrugge:2017fcc ; Feldbrugge:2017mbc ; Matsui:2024bfn (see Lehners:2023yrj for review on the no-boundary proposal). These studies motivated us to investigate whether NBC or RBC can lead to stable behaviour of perturbations, also at the same time highlighting that Dirichlet BC is perhaps not the right boundary condition for gravity Krishnan:2016mcj ; Krishnan:2017bte .
@@
Explorations with the Neumann boundary condition (NBC) have shown that gravitational perturbations are well-behaved in the no-boundary proposal of the Universe DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; Narain:2022msz ; Mondal:2023cxx ; Ailiga:2024nkz . In the case of mini-superspace approximation with metric having ×$\mathbb{R}\times\mathbb{S}^{3}$ it is possible to do the computation of the path-integral exactly Narain:2022msz . These exact computations showed some interesting features where it was seen that our Universe undergoes signature changes from Euclidean to Lorentzian Narain:2022msz . Studies involving RBCs in Einstein-Hilbert gravity were done in DiTucci:2019dji ; Mondal:2022gyp ; DiTucci:2019bui under the WKB approximation using Picard-Lefschetz methods. In all these works it was noticed that perturbations were well-behaved. Exact computations imposing RBCs at the initial boundary have been done in Ailiga:2023wzl . Effects of Gauss-Bonnet addition have been thoroughly investigated in Narain:2021bff ; Narain:2022msz ; Ailiga:2023wzl .
@@
This review also focuses on the impact of having Gauss-Bonnet addition in the gravitational action and the role it plays at the boundaries. This is crucial as the path-integral is sensitive to the boundary conditions. Retaining the consistency of the variational problem, the boundary condition is appropriately chosen. This still doesn’t guarantee that the solutions respecting these will be a stable configuration. But it is crucial to work those boundary conditions which lead to stable perturbations.
@@
Such path-integrals need to be analyzed appropriately in a systematic manner in the arena of complex analysis. Earlier studies involving complex analysis methods to understand Euclidean gravitational path integral was done in Hawking:1981gb ; Hartle:1983ai . Boundary effects studies in the Euclidean quantum cosmology was done in tunnelling proposal made in Vilenkin:1982de ; Vilenkin:1983xq ; Vilenkin:1984wp and no-boundary proposal made in Hawking:1981gb ; Hartle:1983ai ; Hawking:1983hj . Later works investigating the contour choice in the path-integral involving the usage of complex analysis methods were also done in Halliwell:1988ik ; Halliwell:1989dy ; Halliwell:1990qr . More recent work Kontsevich:2021dmb on the ‘allowability’ criterion proposes a simple method to determine the allowable complex metric on which quantum field theories can be consistently defined. It is, however, yet to be understood and established whether this criterion is necessary or sufficient (see also recent works Witten:2021nzp and Lehners:2021mah ). Our approach avoids the standard process of Wick rotation used in the Euclidean approach and aims to directly tackle the Lorentzian gravitational path-integral using the Picard-Lefschetz (PL) method.
@@
Picard-Lefschetz (PL) methodology generalizes the notion of standard Wick-rotation by suitably picking a contour along which oscillatory Lorentzian integrals can be performed in a reliable manner. It provides a framework that considers all the saddles of the path-integral and systematically picks only the ones which eventually lead to convergent path-integral. The original integration contour gets appropriately deformed into the integrals over the steepest descent contours passing through such saddles. Such contours termed Lefschetz thimbles together form the generalized Wick-rotated contour. In the Lorentzian quantum cosmology studies, these methods were utilized to understand the nature of the path integral Feldbrugge:2017kzv ; Feldbrugge:2017fcc ; Feldbrugge:2017mbc ; Vilenkin:2018dch ; Vilenkin:2018oja ; Rajeev:2021xit and the effects of the boundary conditions DiTucci:2019dji ; DiTucci:2019bui ; Narain:2021bff ; Ailiga:2024nkz ; Lehners:2021jmv ; Narain:2022msz ; Ailiga:2023wzl
@@
The plan of this article is as follows: In section I, we present the introduction and discuss the motivation for studying the gravitational path integral. Section II discusses the mini-superspace approximation and outlines the surface terms required for the consistent variational problem. In section III, we compute the exact transition amplitude for Einstein-Gauss-Bonnet gravity with Neumann/Robin boundary conditions. We analyze the semiclassical behaviour along with the significance of Gauss-Bonnet in the section IV. In section IV.3, we discuss the significance of the Robin boundary condition. A brief overview of the Picard-Lefschetz technique is provided in the section V, followed by its application in the section VI to compute the saddle point results and verify their consistency with the results obtained from the exact analysis. Section VII outlines the connection between the path integral and canonical approaches. Finally, the article concludes with a summary and outlook in Section VIII.